OpenStudy (anonymous):

What is the solution of the system? Use substitution. x = –2y + 10 –6y = 6x – 6 z = –3x + 10y A. x = 9, y = –8, z = 114 B. x = –8, y = 9, z = 114 C. x = –8, y = 114, z = 9 D. x = 114, y = 9, z = –8

OpenStudy (anonymous):

@.Sam.

sam (.sam.):

When x = –2y + 10 , From –6y = 6x – 6 –6y = 6( –2y + 10 ) – 6 Solve for y

OpenStudy (anonymous):

? I such at this

OpenStudy (anonymous):

so its C

sam (.sam.):

\[–6y = 6( –2y + 10 ) – 6 \\ \\ -6y=-12y+60-6 \\ \\ 6y=54 \\ \\ y=9\] Get it?

OpenStudy (anonymous):

no

sam (.sam.):

You distribute 6 into brackets \[–6y = 6( –2y + 10 ) – 6\] \[-6y=-12y+60-6\] Add 12y to both sides, 6y=54 Divide both sides by 6 y=9

OpenStudy (anonymous):

both sides of what

OpenStudy (anonymous):

I don't understand this at all and slopes forget it

sam (.sam.):